OpenStudy (tbnrfrags):

The average (arithmetic mean) of 5 numbers is 13. If the average of 4 of these numbers is 10, what is the fifth number?

OpenStudy (anonymous):

We know that five numbers, (x+y+v+z+w)/5 = 13, and that four numbers, (x+y+v+z)/4 = 10. If we say that x+y+v+z = a, then we have two solveable equations, wouldn't you say?

OpenStudy (anonymous):

is it 25..

OpenStudy (tbnrfrags):

yeah im failing this class so i couldnt tell you haha

OpenStudy (anonymous):

You just have to solve for a, then sub it into the second equation to find w. \[\frac{ a }{ 4} = 10.........(1)\] \[\frac{ a+w }{ 5 }.......(2)\]

OpenStudy (anonymous):

Sorry, that second equation should have an = 13 at the end of it

OpenStudy (mathmath333):

\(\large\tt \color{black}{take~~the~~five~~numbers~~as~~(x-3a),(x-a)(x),(x+a)(x+3a)}\)

OpenStudy (cj49):

avg of 5 @13 =65 avg of 4 @10=40 therefore,65-40=25

OpenStudy (cj49):

5th no is 25

OpenStudy (mathmath333):

nice..

OpenStudy (tbnrfrags):

ty all<3